Find the vertex and the line of symmetry of the quadratic function.
F(x)=-2x^2+2x+1
The x coordinate of the vertex is ____. (Type a simplified fraction.)
The y coordinate of the vertex is ____. (Type a simplified fraction.)
The equation line of symmetry is ____. (Type a simplified fraction.)

Answer by rfer(16322)

x co=1/2
y co=3/2
axis x=-b/2a
x=-2/-4
x=1/2